OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 9x2 - 18x - 45 = 0

OpenStudy (anonymous):

9*(x-1)^2-54=0 so now add 54 9*(x-1)^2=54 Div by 9 (x-1)^2=54/9

OpenStudy (anonymous):

(x – 9)2 = 15 <<<--- Would this be the answer?

OpenStudy (anonymous):

(x-9)^2=6

OpenStudy (anonymous):

I'm sorry, it's not listed on the answers; this is like college work, and I'm still in High School.

OpenStudy (anonymous):

it's HS :)

OpenStudy (anonymous):

\[9x ^{2}-18x-45=0\] \[x ^{2}-2x-5=0\] \[x ^{2}-2x+1-6=0\] \[(x-1)^{2}-6=0\] \[(x-1)^{2}=6\]